# PatentScore: Multi-dimensional Evaluation of LLM-Generated Patent Claims

### v1

@@ Abstract
Natural language generation (NLG) metrics play a central role in evaluating generated texts, but are not well suited for the structural and legal characteristics of patent documents. Large language models (LLMs) offer strong potential in automating patent generation, yet research on evaluating LLM-generated patents remains limited, especially in evaluating the generation quality of patent claims, which are central to defining the scope of protection. Effective claim evaluation requires addressing legal validity, technical accuracy, and structural compliance. To address this gap, we introduce PatentScore, a multi-dimensional evaluation framework for assessing LLM-generated patent claims. PatentScore incorporates: (1) hierarchical decomposition for claim analysis; (2) domain-specific validation patterns based on legal and technical standards; and (3) scoring across structural, semantic, and legal dimensions. Unlike general-purpose NLG metrics, PatentScore reflects patent-specific constraints and document structures, enabling evaluation beyond surface similarity. We evaluate 400 GPT-4o-mini generated Claim 1s and report a Pearson correlation of $r=0.819$ with expert annotations, outperforming existing NLG metrics. Furthermore, we conduct additional evaluations using open models such as Claude-3.5-Haiku and Gemini-1.5-flash, all of which show strong correlations with expert judgments, confirming the robustness and generalizability of our framework.
@@ 1 Introduction
Patents involve rapidly increasing, significant technological, economic, and social impacts. More than 3 million patent applications were filed globally in 2021 (WIPO, 2022), representing millions of dollars in the intellectual property (IP) market (OECD, 2021). However, pattern drafting, reviewing, and filing are not only time and financially costly but also technically challenging due to their complex structures, legal constraints, and technical specifications (Surden, 2019). Drafting a patent application could cost 30-40 hours of work by patent attorneys and technical experts, costing USD\$5,000 to \$15,000 (WIPO, 2022). Therefore, automating patent generation and examination has a growing global urgency and priority (Ciccatelli, 2019), where AI, specifically natural language generation (NLG) and large language models (LLMs), play essential roles (Aristodemou and Tietze, 2018). This demand grows rapidly over time, as illustrated in Figure 1 of patent applications and registrations in Australia from 2011 to 2021. Widely required in the global patent domain, further emphasizing the importance of patent automation (Fujii, 2007).
Figure 1: Patent applications and registrations in Australia in 2011-2021 (IP Australia, 2022).
This trend of increasing patent activities is even more pronounced in major economies. China and the United States lead global patent applications (WIPO, 2022). The high application volume at these offices has led to increasing backlogs and processing times, emphasizing the need for AI-assisted patent automation (Lee and Hsiang, 2020). Patent automation has the potential to revolutionize the patent filing process, significantly reducing the time, cost, and human effort involved, and significantly enhancing the efficiency of patent filing and innovation processes. (Ebrahim, 2016).
A specific area concerned in this paper is the automated evaluation of LLM-generated patent Claim 1, which is the most important section in patent applications. Claim 1 defines comprehensive but precise legal and technical scopes of IP value and protection and determines critical legal boundaries of patent rights (Merges, 1997). It is essential to ensure legal and technical accuracy when evaluating AI-generated patent claims. Errors or ambiguities in Claim 1 can lead to litigation, invalidation, or reduced protection, making precise evaluation of generated patent claims critically important (Bessen and Meurer, 2008). Addressing these unique requirements and challenges demands specialized evaluation frameworks for LLM-generated patent claims.
However, this demand cannot be met by existing NLG-based evaluation metrics originally designed for general text evaluation, which focus primarily on lexical and semantic similarities (Papineni et al., 2002; Lin, 2004). For example, BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) primarily evaluate n-gram overlap, and BERTScore (Zhang et al., 2019) focuses on contextual embeddings for semantic similarity. They inherently lack the mechanisms to address critical patent-specific requirements such as legal compliance and technical consistency, and fail to address patent-specific requirements such as structural coherence, claim dependency, antecedent basis, and legal compliance, which are crucial for evaluating patent claims (Merges et al., 1997). In addition, while LLMs have demonstrated substantial progress in transforming many language tasks such as question-answering and machine vision Brown et al. (2020), their applications in generating and evaluating patent claims are unexplored.
Accordingly, we address these limitations by integrating claim structural validation and claim-specific legal criteria into a multi-dimensional claim evaluation framework PatentScore, significantly enhancing claim evaluation accuracy and reliability. Compared with NLG metrics shown in Figure 2 (a), which only analyzes text-level similarity without considering their hierarchical structures or legal specifications, PatentScore shown in Figure 2 (b) systematically analyzes hierarchical claim structures by decomposing claim structures and integrates patent-specific legal validation patterns into effectively evaluating aspects like claim dependency and ambiguous scope. These ensure legal and technical accuracy of generated claims, thus significantly enhancing the reliability of patent claim evaluation.
Figure 2: The comparison between NLG metric and PatentScore.
Illustrated on GPT-4o-mini-generated patent claims, PatentScore provides an effective pattern generation evaluation framework that extends the Chain-of-Thought (Wei et al., 2022) to structural, legal, and semantic aspects of patent claims. It develops and integrates the following three important aspects into LLM-generated claim evaluation:
  • Patent-structured analysis architecture: A novel framework that systematically analyzes patent claims through hierarchical decomposition, validating structural and technical compliance. It incorporates patent-specific requirements such as claim structure and scope verification, essential for legal validity yet overlooked by existing NLG metrics.
  • Patent-specific quality metrics: A comprehensive evaluation introducing new metrics for legal and technical accuracy, leveraging semantic and terminological consistency in patents.
  • Verifiable evaluation protocol: A systematic validation demonstrating a strong correlation with expert judgments (r = 0.819) through a rigorous evaluation at the component level of seven key patent-oriented factors.
Unlike prior studies that primarily focused on novelty assessment or prior art retrieval, this paper addresses a distinct yet critical problem: evaluating the intrinsic quality of individual claims generated by LLMs. As the legal foundation of patent protection, claims demand structural, technical, and legal precision. To this end, we propose PatentScore, a multi-dimensional evaluation framework, and conduct large-scale experiments on claims generated by GPT-4o-mini. The results show a strong correlation with expert judgments ($r=0.819$) and demonstrate that PatentScore significantly outperforms existing metrics such as BLEU, ROUGE, GPTScore, and cosine similarity. PatentScore provides a reliable and generalizable framework for evaluating AI-generated patent claims, serving as a practical foundation for building LLM-based pipelines for automated patent drafting and assessment.
2 Background & Related Work
@@ 1. An extendible trailer tongue comprising:
Table 1: An example of Claim 1 from the patent titled "Multiple component headgear system" (Siprut, 2000).
Patents are specialized legal-technical documents that protect IP rights. Among the various components of a patent, claims define the legal boundaries of IP protection, with Claim 1 serving as the cornerstone. Claim 1 establishes the broadest scope of protection and directly impacts the patent’s validity and enforceability (WIPO, 2023; Merges et al., 1997).
Unlike general technical documents, patent claims require adherence to strict structural formats, standardized terminology, and legal rigor. Table 1 illustrates a well-formed Claim 1, which demonstrates these key requirements:
  • Structural format: Compliance with strict formatting requirements.
  • Technical clarity: Precise definition of technical elements and relationships.
  • Terminology consistency: Use of standardized patent-specific language.
@@ 2.2 Patent Generation and Evaluation
Automating patent generation and evaluation has become an emerging agenda, yet remains underexplored. Large language models (LLMs) have recently been applied to generate technical and legal content Brown et al. (2020), showing promise in zero-shot reasoning Kojima et al. (2023) and complex instruction following Li et al. (2024). However, most existing studies focus on general text generation, without addressing the strict structural and legal requirements of patent claims Surden (2019). Patent evaluation involves both linguistic and domain-specific dimensions. Traditionally, especially Claim 1,are reviewed manually by examiners and attorneys, which is time-consuming and inconsistent. As patent applications increase globally, automated evaluation systems must assess technical correctness, legal soundness, and structural compliance simultaneously (Surden, 2019).
Yet, evaluating LLM-generated patent claims is particularly challenging due to their multi-faceted nature. Minor errors in structure, terminology, or legal scope can result in litigation or invalidation. Existing evaluation metrics such as BLEU, ROUGE, and BERTScore Sellam et al. (2020), primarily assess linguistic similarity but fail to capture critical patent-specific elements like claim scope, legal consistency, and structural validity.
Recent LLM evaluation frameworks, including PromptScore, HolisticEval, and LMdiff (Gao et al., 2024; Wang et al., 2023; Liu et al., 2024), provide task-specific coherence assessments. However, they do not account for the legal and structural constraints inherent to patents. PatentScore addresses this gap by incorporating patent-specific evaluation dimensions into its framework.
@@ 2.3 NLG Metrics for Patent Evaluation
NLG metrics dominate patent generation evaluation, we review them specifically. First, traditional n-gram based metrics like BLEU and ROUGE measure lexical overlap between generated and reference texts. However, these lexical similarities overlook rich semantic relationships and domain specialties Sellam et al. (2020), which are crucial particularly for specialized contexts like patent claims with legal and technical precision Surden (2019). Then, metrics based on deep models like Transformer enable more nuanced evaluation. For example, BERTScore utilizes contextual embeddings to assess semantic similarity, while BLEURT Sellam et al. (2020) combines BERT’s representational power with task-specific optimization. GPTScore Fu et al. (2023) further leverages LLMs to evaluate semantic coherence, logical consistency, and contextual appropriateness, making it promising for legal and technical content evaluation.
Despite these advancements, the NLG/LLM evaluation metrics struggle to address unique challenges posed by patent documents with strong legal features and requirements. Unlike general text, patent claims require rigorous structural compliance, precise legal terminology, and domain-specific validation. Existing metrics fall short in assessing critical elements such as claim scope definition, antecedent basis establishment, and hierarchical relationships between claims. This highlights the necessity for specialized evaluation frameworks capable of addressing both linguistic and patent-specific requirements comprehensively.
@@ 3 PatentScore for Patent Claim Evaluation
Figure 3: The PatentScore Framework.
3.1 The PatentScore Framework for multi-dimensional patent claim evaluation.
PatentScore is a systematic evaluation framework for patent claims that integrates structural, legal, and semantic analyses in a decomposed manner. Figure 3 illustrates how these components work together to generate overall quality scores while maintaining independence in their evaluations.
First, the structural analysis identifies three key characteristics: claim structure ($M_{CS}$), punctuation pattern ($M_{CP}$), and antecedent basis ($M_{AB}$). Each is scored on a scale of 1-5. Then, the legal analysis examines element referencing ($M_{ER}$), validity/uniqueness ($M_{VU}$), and ambiguous scope ($M_{AS}$), also scored on a scale of 1-5. For semantic analysis, we employ BERTScore ($M_{BS}$), a proven metric in NLG, which provides scores between 0 and 1, later scaled to 1–5 for consistency with other metrics. To systematically evaluate these components, we leverage GPT-4o-mini with carefully designed structural, legal and semantic prompts for them. Each component is assessed through specialized prompts that guide the model to focus on specific aspects below. Detailed prompts are provided in Appendix A.
Structural component
evaluates the technical composition and formatting of claims, following standard patent writing conventions and WIPO patent drafting guidelines WIPO (2023).
Legal component
analyzes compliance with patent law requirements and clarity standards based on USPTO’s MPEP guidelines USPTO (2023) and WIPO patent drafting guidelines WIPO (2023).
Semantic component
employs BERTScore to measure semantic similarity, a proven approach in NLG for assessing textual coherence.
These components apply specific prompts that enable PatentScore to perform targeted evaluation of each metric while maintaining consistency across assessments. PatentScore enables decomposing the complex task of patent claim evaluation into well-defined sub-tasks that can be systematically analyzed using LLMs. To mitigate the inherent variability of GPT model outputs, we perform ten iterations of evaluation for each metric and use the mean score as the final value. To further enhance output consistency, we fix the temperature parameter at 0.3. This configuration ensures more reliable and stable evaluation results. Furthermore, we validate the robustness and generalizability of PatentScore by applying the same evaluation pipeline to multiple LLMs, including GPT-4o-mini, Claude-3.5-Haiku, and Gemini-1.5-flash. Since these evaluation criteria are grounded in internationally recognized guidelines (e.g., WIPO (WIPO, 2023), USPTO (USPTO, 2023)), PatentScore can be readily adapted to other jurisdictions with minimal adjustments.
3.2 Decomposed Patent Evaluation Components
We propose patent evaluation components in structural, legal, and semantic aspects, respectively.
Structural Components consist of claim structure, claim punctuation, and antecedent basis.
Claim Structure ($M_{CS}$): Assesses the preamble, transitional phrases, and elements to ensure structural completeness and clear enforceability.
Claim Punctuation ($M_{CP}$): Focuses on critical punctuation marks (e.g., colons, semicolons, "and," periods) to maintain compliance with patent drafting standards and ensure accurate interpretation.
Antecedent Basis ($M_{AB}$): Evaluates the proper establishment of antecedent basis by checking article consistency in accordance with patent drafting standards.
Legal Components comprise element referencing, validity and uniqueness, and ambiguous scope.
Element Referencing ($M_{ER}$): Verifies the accuracy and consistency of references to technical components, ensuring clear relationships and dependencies among claim elements.
Validity and Uniqueness ($M_{VU}$): Evaluates whether claim elements exhibit technically distinctive functions or expressive originality beyond common domain knowledge, based on prompt-guided assessment.
Ambiguous Scope ($M_{AS}$): Identifies and resolves overly broad or ambiguous language, ensuring the claim scope is precise and legally enforceable.
Semantic Components consist of BERTScore.
BERTScore ($M_{BS}$): Measures semantic similarity and contextual alignment of claims with reference descriptions, ensuring accurate representation of domain-specific terminology.

### v2

@@ Abstract
High-stakes texts such as patent claims, medical records, and technical reports are structurally complex and demand a high degree of reliability and precision. While large language models (LLMs) have recently been applied to automate their generation in high-stakes domains, reliably evaluating such outputs remains a major challenge. Conventional natural language generation (NLG) metrics are effective for generic documents but fail to capture the structural and legal characteristics essential to evaluating complex high-stakes documents. To address this gap, we propose PatentScore, a multi-dimensional evaluation framework specifically designed for one of the most intricate and rigorous domains, patent claims. PatentScore integrates hierarchical decomposition of claim elements, validation patterns grounded in legal and technical standards, and scoring across structural, semantic, and legal dimensions. In experiments on our dataset which consists of 400 Claim1, PatentScore achieved the highest correlation with expert annotations ($r=0.819$), significantly outperforming widely used NLG metrics. This work establishes a new standard for evaluating LLM-generated patent claims, providing a solid foundation for research on patent generation and validation.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in automated text generation across a wide range of natural language processing tasks. However, existing evaluation frameworks remain insufficient for rigorously and systematically assessing their outputs in high-stakes domains. These domains, such as patent claims, legal contracts, and technical reports, involve deeply hierarchical and interdependent structures that cannot be accurately evaluated using surface-level metrics such as sentence-level fluency or shallow contextual similarity. Current evaluation metrics often fail to capture critical domain-specific nuances, including legal constraints, structural dependencies, technical precision, and implicit reasoning, which significantly limit their effectiveness in these contexts. This limitation is especially consequential because it poses risks to both legal validity and practical utility of the generated documents, thereby hindering their adoption in real-world, risk-sensitive applications (Guha et al., 2023).
Figure 1: A comparison of PatentScore and standard NLG metrics in evaluating structural and legal accuracy of patent claims.
Among high-stakes legal and technical documents, patent claims pose particularly significant challenges for evaluation. Because they function as both technical descriptions and legally binding texts, patent claims require precise assessment of structural elements such as antecedent consistency and claim dependency. Existing natural language generation (NLG) evaluation methods struggle in domains where both linguistic precision and legal accountability are critical. Patent claims exemplify these characteristics, as they define the scope of rights granted to an invention. The first claim, in particular, plays a central role in shaping the structure and coverage of the entire claim set. This pivotal role entails substantial socio-economic implications, as it determines the legal and commercial value of the invention (Merges, 1997; Bessen and Meurer, 2009; Surden, 2019).
With the rapid advancement of LLMs, recent studies have explored their potential to generate and evaluate precision-critical legal texts. However, patent claims remain a particularly challenging domain due to their intricate structural and legal constraints (Aristodemou and Tietze, 2018; Lee and Hsiang, 2020). Traditional NLG metrics, such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and BERTScore (Zhang et al., 2020), primarily focus on lexical overlap or contextual similarity, failing to address the structural and legal nuances required for patent claim evaluation.
More recent LLM-based evaluation approaches, such as LLM-as-a-Judge (Zheng et al., 2023), rely on general-purpose models to assess fluency and semantic coherence. However, they lack the domain-specific knowledge required to evaluate structural constraints, including claim dependency and legal clarity, that are critical to patent claims. These shortcomings limit the applicability of such metrics in high-stakes legal contexts where precision is paramount (Zuo et al., 2024). For instance, inconsistent antecedent references can render a patent legally ambiguous, while punctuation errors may distort its scope of protection (Merges, 1997).
To address these limitations, we propose PatentScore, a multi-dimensional evaluation framework for patent claims. This framework evaluates structural and legal elements such as antecedent consistency, claim dependency, and legal clarity. Our approach decomposes claims into constituent parts and applies targeted metrics, enabling reliable and consistent assessment of LLM-generated claims. This systematic evaluation significantly enhances the quality of claim evaluation compared to existing metrics. Our objective is to develop a domain-specific metric for patent claims while establishing a generalizable, structure-aware evaluation framework applicable to high-precision documents, such as contract clauses, policy reports, and regulatory compliance in clinical documentation.
The key contributions of PatentScore are summarized as follows:
  • Claim-Structured Evaluation Framework. We introduce an architecture that decomposes a claim into structural components and quantifies each, covering claim structure, element linkage, antecedent basis, and claim dependency per official examination guidelines.
  • Expert-Aligned Validation. We compare PatentScore’s evaluation of LLM-generated claims against expert ratings, achieving a strong correlation ($r=0.819$) and outperforming standard NLG metrics in capturing structural and legal precision.
  • Open Evaluation Benchmark. We release 400 LLM-generated claims with expert annotations, creating the first public benchmark for evaluating LLM-generated patent texts.
Although our experiments focus on patent claims, PatentScore is model-agnostic. Its structural and legal principles extend to other high-precision texts such as contract clauses, policy reports, clinical documentation, and regulatory filings. By providing a robust framework for evaluating structured texts, our work lays the foundation for establishing reliable evaluation protocols essential for the responsible use of LLMs in high-stakes domains.
2 Background and Related Work
@@ 1. An extendible trailer tongue comprising:
Table 1: A Brief summary example of Claim 1 from the patent titled “Multiple component headgear system” (Siprut, 2000).
Patents are specialized legal-technical documents designed to protect intellectual property (IP) rights. Among their components, claims play a central role, as they define the precise legal boundaries of protection granted to an invention. In particular, Claim 1, which is the first claim, typically serves as the cornerstone of the claim set. It establishes the broadest scope of protection and strongly influences both the validity and enforceability of the patent as a whole (World Intellectual Property Organization, 2022; Merges, 1997). The breadth of Claim 1 not only constrains subsequent dependent claims but also directly affects the legal and commercial value of the invention.
Unlike general technical documents, patent claims must strictly adhere to structural, linguistic, and legal requirements. They follow a standardized format mandated by patent offices, employ precise domain-specific terminology, and are drafted with enforceability in mind. Table 1 illustrates an example of a well-structured Claim 1, which demonstrates these key requirements:
  • Structural format: Compliance with strict grammatical and formatting conventions, including numbered clauses, transitional phrases (e.g., “comprising”), and hierarchical listing of elements. Deviations can render a claim unclear or legally unenforceable.
  • Technical clarity: Precise definition of technical element and its relationships, ensuring that the scope of the invention is unambiguous to examiners, practitioners, and courts.
  • Terminology consistency: Reliance on standardized patent-specific language and consistent referencing of terms, minimizing ambiguity and potential disputes.
Together, these interrelated requirements ensure that Claim 1 functions simultaneously as a detailed technical specification and as a legally binding definition of the scope of the invention. Drafting Claim 1, therefore, requires a delicate and deliberate balance: it must be broad enough to secure meaningful protection and commercial value, while being sufficiently precise and consistent to withstand rigorous legal scrutiny.
@@ 2.2 Patent Generation and Evaluation
Automating patent generation and evaluation has recently emerged as an important research agenda, yet it remains relatively underexplored. LLMs have been increasingly applied to generate technical and legal content (Brown et al., 2020), showing promise in zero-shot reasoning (Kojima et al., 2022) and complex instruction following (Lai et al., 2024). However, most existing work has focused on general text generation, without addressing the strict structural and legal constraints unique to patent claims (Surden, 2019).
Patent evaluation inherently involves both linguistic and domain-specific dimensions. Traditionally, examiners and attorneys manually review claims, particularly Claim 1, which is labor-intensive and often inconsistent. With the rapid global increase in patent applications, automated evaluation systems must be able to simultaneously assess technical correctness, legal soundness, and structural compliance (Surden, 2019).
Nevertheless, evaluating LLM-generated patent claims remains a highly challenging and technically demanding task. Even minor errors in structure, terminology, or legal scope can result in costly litigation or eventual invalidation of the patent. Existing NLG metrics tend to emphasize surface-level linguistic similarity or superficial fluency, but they fail to capture critical patent-specific aspects such as claim scope, legal consistency, structural validity, and the intricate interdependence of claim elements. This limitation strongly underscores the necessity for domain-specific evaluation frameworks that can comprehensively and reliably address the unique requirements of patent claims.
@@ 2.3 NLG Metrics for Patent Evaluation
Traditionally, the evaluation of patent texts has relied on metrics using n-grams such as BLEU and ROUGE, which measure lexical overlap between generated and reference texts. However, these approaches overlook semantic relationships and domain-specific nuances, and thus are insufficient for patent claims that demand high levels of legal and technical precision. With the introduction of Transformer-based models, more nuanced evaluation methods became possible. For example, BERTScore leverages contextual embeddings to assess semantic similarity (Zhang et al., 2020), while BLEURT combines BERT’s representational power with task-specific optimization (Sellam et al., 2020).
More recently, LLM-based evaluation frameworks have emerged, directly employing large language models as evaluators. GPTScore exploits the reasoning capability of LLMs to assess semantic coherence, logical consistency, and contextual appropriateness (Fu et al., 2024). Other studies further highlight the potential of LLM-as-a-Judge (Zheng et al., 2023), showing that LLM evaluators can often align more closely with human judgment compared to traditional automatic metrics.
However, these approaches still fail to adequately capture the strict legal and structural constraints inherent in patent claims. Critical aspects such as claim dependency, antecedent basis, and legal clarity remain largely beyond the practical scope of existing evaluation metrics. Unlike general text, patent claims require rigorous structural compliance, precise legal terminology, and systematic domain-specific validation. This persistent gap underscores the urgent need for specialized evaluation frameworks that can comprehensively address both linguistic quality and patent-specific requirements. PatentScore is explicitly designed to fill this void by holistically integrating structural, semantic, and legal dimensions into a unified framework tailored for patent claims.
Figure 2: PatentScore Evaluation Framework: decomposition into structural, legal, and semantic dimensions, normalization of metrics, and weighted aggregation for final scoring.
@@ 3 PatentScore for Patent Claim Evaluation
3.1 Evaluation Dimensions of PatentScore
PatentScore is a multi-dimensional evaluation framework that assesses the quality of patent claims by decomposing them into three independent aspects: structural, legal, and semantic. Each dimension is grounded in domain-specific standards, particularly the WIPO Patent Drafting Manual World Intellectual Property Organization (2022) and the USPTO USPTO (2023), and is operationalized through distinct component metrics. These dimensions operate independently, allowing for a detailed diagnosis of various characteristics of a patent claim.
(a) Structural Dimension.
The structural dimension evaluates the syntactic organization and formal composition of patent claims. It focuses on whether the claim conforms to internationally accepted formatting rules that promote legal clarity and technical completeness. This dimension comprises the following three components:
  • Claim Structure ($M_{CS}$): Checks whether the claim follows the canonical structure, including a preamble, constituent elements, and functional linkages, and logical coherence between interconnected parts.
  • Claim Punctuation ($M_{CP}$): Verifies proper punctuation use (e.g., commas, semicolons) that affect precise clause separation and legal interpretation.
  • Antecedent Basis ($M_{AB}$): Assesses the proper usage of definite noun phrases (e.g., “the module”, “said component”) that must be supported by clear prior mention.
(b) Legal Dimension.
The legal dimension assesses whether the claim satisfies enforceability criteria under patent law, referencing legal doctrine from the USPTO’s MPEP and WIPO standards. Its core metrics include:
  • Element Referencing ($M_{ER}$): Measures proper dependency and referencing in independent and dependent claims.
  • Validity and Uniqueness ($M_{VU}$): Evaluates novelty, non-triviality, and absence of internal contradiction.
  • Ambiguous Scope ($M_{AS}$): Identifies vague or overbroad terms that could undermine legal clarity or enforceability.
(c) Semantic Dimension.
The semantic dimension evaluates the extent to which a generated claim preserves the intended meaning and technical context of the original. We adopt BERTScore ($M_{BS}$), which measures semantic similarity using contextual embeddings, due to its robustness in capturing meaning beyond lexical overlap. To ensure compatibility with other dimensions, BERTScore values in $[0,1]$ are linearly rescaled to a $[1,5]$ range.
This decomposition enables fine-grained diagnostics of claim quality, enhancing both the reliability and interpretability of LLM-based assessments.
3.2 Implementation of PatentScore
To operationalize the three evaluation dimensions, PatentScore employs an LLM-based prompt evaluation scheme. For each metric, a dedicated prompt is designed to guide the model’s attention to the corresponding evaluation criterion. GPT-4o-mini is used as the primary evaluator, and ten independent evaluations are conducted for each metric, with the average score adopted as the final result. This design reduces output variance and ensures stability in scoring.
Different evaluation strategies are applied depending on the nature of each metric. For structural and legal metrics, explicit prompts are used to focus the model on the core evaluation concepts. In contrast, the semantic metric is assessed using BERTScore, which is computed independently. The original $[0,1]$ range of BERTScore values is linearly rescaled to a $[1,5]$ range for consistency with other scores. All prompt templates are provided in detail in Appendix A.
The proposed implementation is model-agnostic and can be applied to various LLMs, including GPT-4o-mini and others. Since the evaluation criteria are grounded in internationally recognized patent drafting guidelines (WIPO, USPTO), the framework ensures generalizability and applicability across different contexts.